THE VIBRATIONAL ASSIGNMENT OF NO_{3} IN THE GROUND ELECTRONIC STATE

\underline{EIZI HIROTA}, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, Japan; KENTAROU KAWAGUCHI, Department of Chemistry, Okayama University, Tsushima-naka 3-1-1, Okayama 700-8530, Japan; TAKASHI ISHIWATA, Department of Information Machines and Interfaces, Faculty of Information Sciences, Hiroshima City University, Ozukahigashi, Asaminami, Hiroshima 731-3194, Japan; IKUZO TANAKA, Tokyo Institute of Technology, 2-12-1, Ookayama, Meguro, Tokyo 152-8550, Japan.

We have previously analyzed a band of the NO_{3} radical observed at 1492cm^{−1}and have established that the band was of E type, thus being assigned to the degenerate N-O stretching, ν_{3} ^{a}. We have explained several anomalies noticed in the band in terms of a vibronic interaction model ^{b}. Stanton ^{c} recently proposed an alternative assignment ν_{1} + ν_{4} for this band, primarily based on an ab initio calculated potential. In order to establish the vibrational assignment of the band, we applied the vibronic interaction modelc to the ^{14}NO_{3}-^{15}NO_{3} isotope shift and calculated it to be 16 cm^{−1}in reasonable agreement with the observed value of 20 cm^{−1}^{d} when the band is ν_{3}, whereas we obtained a very small value (about 1 cm^{−1}or less) for Stanton’s assignment, at variance with the observed data. We have also scanned the region from 700 up to 1400 cm^{−1}to detect the ν_{3} band predicted by Stanton at 994 cm^{−1}^{e}, by using a Fourier transform spectrometer. Although some part of this region was covered by strong absorption of the precursor HNO_{3}, we confirmed that there were no bands observed between 925 and 1277 cm^{−1}that were more than 1/10 as intense as the 1492 cm^{−1}band.

aT. Ishiwata etal., J. Chem. Phys. 82, 2196 (1985), K. Kawaguchi etal., J. Chem. Phys. 93, 951 (1990).

bE. Hirota etal., J. Chem. Phys. 95, 771 (1991).

cF. Stanton, J. Chem. Phys. 126, 134309 (2007).

dT. Ishiwata etal., J. Chem. Phys. 82, 2196 (1985).

eF. Stanton, J. Chem. Phys. 126, 134309 (2007).